EVA Report – February 11th

Crew 221 EVA # 10 Report 11 Feb 2020

EVA # 10

Author: Luis Monge (Engineer)

Purpose of EVA: We will be performing the deployment of a dipole antenna and document the whole process. The antenna and required implements will be carried in the rovers and assembled close to them at the destination.

Start time: 10:00 am

End time: 11:30 pm

Narrative: Today we conducted the first of the two EVA for antenna deployment exercises. L.Monge, and G.Alloitabi went out to drive towards Cowboy Corner with the aim to practice long wave antenna installation. The Martian weather was sunny and chilly during the EVA. L.Monge took the lead and drove directly to the spot, once there proceeded to assemble the antenna.

The procedure was documented with video and pictures, the antenna was disassembled the components stored in the rovers and the crew headed back home on time.

Destination: Cowboy Corner

Coordinates (use UTM NAD27 CONUS): 4252600/518770 (Cowboy Corner)

EVA Participants: Luis, Ghanim

Road(s) and routes per MDRS Map: Road(s) and routes per MDRS Map: Park along Cow Dung Road (0110) to stop at Cowboy Corner

Mode of travel: Driving and Walking

Vehicles you will be using (If applicable): Curiosity and Spirit

EVA Report – February 10th

Crew 221 EVA Report 10FEB-2020

EVA # 9

Author: Jessica Todd (GreenHab Officer)

Purpose of EVA: We performed a duplicate EVA of this morning’s EVA, trekking over to the North Ridge and up along Hab Ridge to collect 8 samples for microbiology studies. We attempted to follow instructions based on landmarks located by EVA team 1, and follow a designated path from the terrain map. We will be comparing the results of the efficiency, traversing and science of the two EVAs.

Start time: 2:50 pm

End time: 5:09 pm

Narrative: Our EVA team set off from the Hab heading north towards North Ridge. We were aiming for the metal pole marker, indicating the Hab Ridge Rd. We got a little lost at the marker, trying to locate the path, but after a bit of exploring we found the ATV trail that indicated the start of the path up the ridge. We hiked up Hab Ridge, taking samples on the way up, however progress was slow due to one of the EVA team members having a sore knee. Once we reached the top of the ridge, the wind started to pick up, so we turned around and headed back down, taking some more samples on the descent. On the way back to the Hab, one of our radios ran out of charge, so we communicated with hand signals to navigate the path back. We had caught a mouse in the Hab earlier that day, and so with the permission of the mission support team, two EVA members retrieved the trapped mouse from the airlock, swapped out the faulty comms, and headed back out on two rovers. The team member with an injured knee remained behind and finished up their EVA. We drove the two rovers out to Cowboy Corner, let loose Kim the Mouse and returned to the Hab.

Destination: North end of Hab Ridge (EVA). Cowboy Corner (Mouse drop off)

Coordinates (use UTM NAD27 CONUS): 0517500/4251800 (Hab Ridge), 0518700/4252650 (Cowboy Corner)

Participants: Jessica Todd (GreenHab Officer), Hemani Kalucha (Journalist), Rawan Alshammari (HSO)

Road(s) and routes per MDRS Map: Walked north towards North Ridge, turned onto Hab Ridge Rd (1103) and up to Hab Ridge. After returning to the Hab, we picked up Spirit and Curiosity to drop off Kim the Mouse at Cowboy Corner north of Galileo Rd.

Mode of travel: Driving and Walking

EVA Report – February 10th

Crew 221 EVA # 8 Report 10 Feb 2020

EVA # 8

Author: Luis Monge (Engineer)

Purpose of EVA: We will be performing a similar EVA both in the morning and the afternoon to study EVA operations, team dynamics and scientific productivity when adapting to variable terrain. Team 1 in the morning was tasked with collecting a set number of scientific samples and make field geography observations over a specified area, while attempting to follow a preplanned route (based off terrain maps).

Start time: 10:00 am

End time: 12:30 pm

Narrative: Today we conducted the first of the two walking EVA. L.Monge, and M.Grulich went out to walk towards North Ridge with the aim to collect samples mainly for microbiology projects. The Martian weather was sunny and hot during the EVA. L.Monge took the lead and some problems were faced in order to find the route at first. The team stopping on the way to check the maps and road conditions and climbed two small hills to determine the location of the road.

Once on the road, Crew Commander M.Grulich was looking for geology traces where microbes could be found. Samples at eight different locations were collected. Including cotton rubbing of a rock that seemed to have microbial colonies on top of it.

Destination: North end of Hab Ridge

Coordinates (use UTM NAD27 CONUS): 0517500/4251800

EVA Participants: Maria, Luis, Ghanim

Road(s) and routes per MDRS Map: Walk up Hab Ridge Rd (1103), through Schubert Pass to the north end of Hab Ridge

Mode of travel: Walking

Vehicles you will be using (If applicable): NA

EVA Report – February 08th

Crew 221 EVA Report 08FEB-2020

EVA # 7

Author: Jessica Todd (GreenHab Officer)

Purpose of EVA: The primary purpose was to determine test sites for our Mars rover. EVA#6 photographed some interesting areas of potential interest in Candor Chasma, which we wanted to further investigate. Our secondary purpose was to collect additional geological samples from the stratification layers of the chasma.

Start time: 10:03 am

End time: 1:00 pm

Narrative: Today we decided to return to Candor Chasma, based on some interesting photos taken by EVA #6. We were looking for challenging terrain in which to test our rover, as well as sites of geological interest in the canyon. Driving along Cow Dung Road, we at first missed the Stream Bed Connector trail, as it is not marked and the recent weather has washed away part of the start of the trail. We realized once we hit Galileo Rd, turned our rovers around and headed back towards Stream Bed Connector, passing some very confused Martian tourists on the way. Confirming the start of Stream Bed Connector with our GPS and terrain map, we headed off towards the Chasma. EVA suit #7’s fan started to malfunction at the start of the EVA, delaying our start slightly, but was fixed with a quick power cycle. The walk along Stream Bed Connector took more than 20 minutes, as the sand was very moist and difficult to walk on, particularly with our heavy packs and equipment. We first took some time to explore along the lip of the Chasma, photographing the various rock strata and terrain types. It took some time to find an appropriate location to make our way down into the Chasma, and one of the EVA crew felt the terrain was too challenging for them to make their way down. It took another 20 minutes for the rest of the EVA team to find a suitable path down into the Chasma that all members could navigate. As we were running short on EVA time, the decision was made to focus primarily on collecting photos and locations of rover sites, rather than collection of geological samples. We left the sample equipment at the mouth of the Chasma to be picked up on our return, and explored roughly 1 km of the Chasma, taking photos of various rock overhangs, caves and terrain too challenging for a human in an EVA suit to explore. This will be perfect for rover exploration, provided the sand on the Chasma floor is not too moist! The weather today was extremely warm compared to previous EVAs (roughly 16 degrees) and we were extremely hot in our EVA suits, so stopped for water on the way back to the rovers, returning a little early due to the hot weather.

Destination: Walking along Stream Bed Connector and Candor Chasma.

Coordinates (use UTM NAD27 CONUS): 0520285/4251221

Participants: Jessica Todd (GreenHab Officer), Ghanim Aloitabi (Astronomer), Rawan Alshammari (HSO)

Road(s) and routes per MDRS Map: Drove along Entrance Road then north along Cow Dung Road (0110). Parked at entrance to Stream Bed Connector and walked along it to Candor Chasma

Mode of travel: Driving and Walking

EVA Report #6 – February 7th

Crew 221 EVA Report 07 Feb 2020

EVA # 6

Author: Luis Monge (Engineer)

Purpose of EVA: Collection of microbial samples at Candor Chasma deeper into the canyon for geobiology experiments being conducted in the science dome. Collection of more regolith-type substrates for the scientific objectives of the ISRU project (growing plants in Martian soil).

Start time: 09:55 am

End time: 12:30 pm

Narrative: Today we planned our second long EVA. L.Monge, H.Kalucha and M.Grulich went out to drive towards Candor Chasma with the aim to collect samples mainly for microbiology projects. The Martian weather is currently very much in our favor and at some point during the EVA all participants complained about the heat. The sun shining high and there was almost no wind. M.Grulich took the lead on spirit to guide the team stopping on the way to check the maps and road conditions. Once near the destination, the EVA crew left the rovers.

Once at Candor Chasma, Crew Commander M.Grulich was looking for geology traces where microbes could be found. Samples at seven different locations were collected, and in one location it was dug to get a sample form at least 40cm deep into the ground.

While taking samples for microbiology, Crew Commander M.Grulich was very careful using extra nitrile gloves on top of the EVA gloves and sterile sample containers and little shovels that have had been cleaned prior to the EVA. On the way back to the rovers a bag of regolith was collected for Greenhab experiments. All objectives of the EVA have been accomplished.

Destination: Candor Chasma.

Coordinates (use UTM NAD27 CONUS): 519750/4251900 (Parking), Stream Bed connector; 520378/4251094

EVA Participants: Maria Grulich (Commander), Luis Monge (Engineer), Hemani Kalucha (Journalist)

Road(s) and routes per MDRS Map: North on Cow Dung Road (0110) to Galileo Road (1104), East along Galileo Road and park the rovers. Walk along Cactus Road towards Candor Chasma.

Mode of travel: Driving and walking

Vehicles used: Spirit and Curiosity

EVA Report – February 5th

EVA # 5

Author: Hemani Kalucha

Purpose of EVA: Fix weather station instruments, Collection of regolith and geological samples to be used as an ISRU growth substrate for some botanical experiments in the GreenHab, and geology and microbiology experiments.

Start time: 2:10 PM

End time: 3:10 PM

Narrative: We traveled up two mounds in the north ridge, and collected microbiology samples for our agar experiments. We followed a new procedure, which the team decided on in the science discussion yesterday. We wore sterile lab gloves on top of our EVA gloves to collect samples. We used sterile cotton sticks to collect the sample, and sealed each sample in an individual new ziplock bag. These are all the measures we could think of to prevent any organic contamination and cross contamination to our samples. We managed to collect four samples, which was on par with our objectives.

Destination: North Ridge

Coordinates (use UTM NAD27 CONUS): (517900,425100)

Participants: Hemani, Rawan

Road(s) and routes per MDRS Map: Hab Ridge Road

Mode of travel: Walking

EVA Report – February 5th

Crew 221 EVA Report 05 Feb 2020

EVA # 1

Author: Maria Grulich (Commander)

Purpose of EVA: Collection of regolith and geological samples to be used as an ISRU growth substrate for some botanical experiments in the GreenHab, and geology and microbiology experiments.

Start time: 09:35 am

End time: 12:30 pm

Narrative: Today we planned our first long EVA. Before going on EVA, crew commander M.Grulich took a sample of the bacteria on the space suits, in particular, the cleaned EVA suit and the suit after coming back to the Hab to see how contaminated the space suits are getting during an EVA. L.Monge, J.Todd and M.Grulich went out to drive towards the Moons with the aim to collect samples for ISRU and microbiology projects. The Martian weather is currently very much in our favor. The sun was smiling at us and there was almost no wind. M.Grulich took the lead on spirit to guide the team towards the Moons. Once near the Moons, the EVA crew left the rovers along Cow Dung Road. Unfortunately, the EVA preparations took one hour longer than expected therefore the first planned stop at Cowboy Corner was canceled. The EVA team did not manage to identify Brahe Highway 1572 therefore the team probably stopped near Factory Butte View and turned left into the field.

Once at the Moons, crew engineer Monge identified the first out of four locations to collect soil for the ISRU project to grow black corn in it. Crew Commander M.Grulich was looking for geology traces where in former times water could have run through to find and cultivate microbes. One location seems to be very promising – the soil showed bubbles which indicates bacteria activities. While taking samples for microbiology, the team was very careful using extra gloves on top of the EVA gloves and sterile sample containers and little shovels that have had been cleaned prior to the EVA. Also, one air sample was taken to see if there are any microbes in the Martian air around the Moon area. J. Todd was also taking geology samples to investigate any traces of salt indicating former water on the planet. All objectives of the EVA have been accomplished. Shortly before returning to the airlock the EVA team, was checking on the weather station installed yesterday and observed that the water was frozen, making it impossible to read out measurements. This was reported to the support team in the habitat to take action in the second EVA of that day.

The EVA team was working efficiently and very well together. A secondary objective of each EVA is to trace the actual taken road and average heart rate. Unfortunately after 20min, the tracking device stopped recording, which will be investigated in further EVAs to analyze and improve future EVA preparation.

Destination: Gray Moon, Beige Moon, The Moon

Coordinates (use UTM NAD27 CONUS): 4254500/516500 (Gray Moon), 4254350/516500 (Beige Moon), 4254800/515700 (The Moon)

Participants: Maria Grulich (Commander), Luis Monge (Crew Engineer), Jessica Todd (GreenHab Officer)

Road(s) and routes per MDRS Map: Cow Dung Road (0110) parked near Factory Butte View walked into the field.

Mode of travel: Driving and Walking

Vehicles used: Curiosity and Spirit

EVA Report – February 4th

Crew 221 EVA Report 04-02-2020

EVA # 1

Author: Jessica Todd (GreenHab Officer)

Purpose of EVA: The primary purpose was to set up several weather experiments. We were affixing two thermometers, a barometer and a rain gauge to the exterior of the Science Dome tunnel. We were also took several photos for outreach.

Start time: 12:06 pm

End time: 12:50 pm

Narrative: Today was our first maintenance EVA. The GreenHab Officer and Crew Astronomer were setting up several weather experiments as part of our goal to incorporate a Martian weather station into our MDRS mission. We spent the morning constructing a barometer and rain gauge, with the goal of affixing these and two thermometers to the exterior of the MDRS to measure weather conditions at our Hab site. We first assessed the GreenHab tunnel as a potential site for the weather experiments, however the entire tunnel is covered by a tarp and therefore rain runoff would confound our measures. We then headed to the Science Dome and found a strip of tunnel close to the Dome without tarp. We attempted to affix the thermometers to the tunnel exterior with cable ties, however the cable ties kept snapping, potentially because of the temperature, or because they have fatigued with age. We managed to affix the barometer and rain gauge to the base of the tunnel, resting on the floor, using pliers to assist in tightening the cable ties, a task we found particular difficult in our spacesuit gloves. We returned to the Hab with the thermometers having
unfortunately not met all of our objectives. We will need to determine a new method for affixing the experiments and a subsequent maintenance EVA will be required to mount the instruments.

Destination: Vicinity of the Hab.

Coordinates (use UTM NAD27 CONUS): 518500/4250050

Participants: Jessica Todd (GreenHab Officer), Ghanim Aloitabi (Astronomer)

Road(s) and routes per MDRS Map: EVA conducted near the MDRS, at the exterior of the Science Dome and GreenHab.

Mode of travel: Walking

EVA Report #2 – February 3rd

Crew 221 EVA Report 03-02-2020

EVA # 2

Author: Jessica Todd (GreenHab Officer)

Purpose of EVA: The primary purpose was exploration of the immediate environment around the Hab for sites of potential geological interest and familiarity with EVA equipment for the crew members. EVA crew members also practiced proper protocols for collection of geological samples.

Start time: 14:05

End time: 15:14

Narrative: This was the first Martian EVA for two of our crew members, with the EVA led by one of our veteran MDRS astronauts. Setup was a little faster than our morning EVA, we were now a well-rehearsed team! We set off at 14:05, headed for Quarry Road along Entrance Road. We cut across country once we reached Quarry Road, headed over to the white outcroppings of Marble Ritual, and took some samples of the white powdery rocks that we will analyze back in the Science Hab. We collected several samples around the Marble Ritual area of varying rock types, practicing our field geology techniques with a rock hammer and shovel, and cataloging the various sample sites. We found some interesting rocks with a greenish hue which we look forward to analyzing back in the lab. Potential for life perhaps? Around 40 minutes into the EVA the wind started to pick up considerably, so our EVA lead made the decision to head back to the Hab early.

Destination: Marble Ritual via Quarry Road.

Coordinates (use UTM NAD27 CONUS): 518750/4250900

Participants: Jessica Todd (GreenHab Officer), Hemani Kalucha (Journalist) and Ghanim Aloitabi (Astronomer)

Road(s) and routes per MDRS Map: Entrance Road and then North on Quarry Road

Mode of travel: Walking

EVA Report #1 – February 3rd

Crew 221 EVA Report 03-02-2020

EVA # 1

Author: Jessica Todd (GreenHab Officer)

Purpose of EVA: The primary purpose was exploration of the immediate environment around the Hab for sites of potential geological interest and familiarity with EVA equipment for the crew members. EVA crew members also practiced proper protocols for collection of geological samples.

Start time: 10:16

End time: 12:01

Narrative: The EVA started 15 past the hour, the crew of three walked to the main road and from there to Marble ritual, two places where visited. A total of seven samples were collected, two at the first spot, one on the way to the second spot and four at the second spot, including three samples from single multiple layer rock wall. Upon return to the base, the samples were stored and will be cataloged.

Destination: Marble Ritual via Quarry Road.

Coordinates (use UTM NAD27 CONUS): 518750/4250900

Participants: Maria Grulich (Commander), Rawan Abdullah (HSO) and Luis Monge (Engineer)

Road(s) and routes per MDRS Map: Entrance Road and then North on Quarry Road

Mode of travel: Walking